OpenStudy (anonymous):

if the product of some number and 5 is increased by 4, the result is 7 times the number. find the number

OpenStudy (anonymous):

5x+4=7x

OpenStudy (anonymous):

2x=4 x=2

OpenStudy (anonymous):

5x+4=7x 7x-5x=4 2x=4 x=4/2 solve :D